Please use the method of the left and right end internal force and moment of each section. Thanks.This is a similar problem of 5.44 in the book (7th edition), however, you are required to draw the shear and bending-moment d
Answer #1

Given dak FBP EL YkN TO EFx=0 RAx=0 E RANT Ray {Fy=0 fry im 16V to MB=0 Ray + Rey = 4+4=8 link -3 X Ray 7 (1x4) +0.504) Ray 응Areas of shear diagram- A to c Suda U) x2 = 2IN.M C to D Srda = UX-2 = -2/N.m D to E Jvda (1)X = -6kn.m Bending moment diagra